OpenStudy (lgg23):

Solve for x: (9/7)log base 2 of x = -9

OpenStudy (anonymous):

First get the log function by itself: \[\frac {7}{9}*\frac {9}{7} *\log_{2}(x)=\frac{-9}{1}*\frac {7}{9}\]

OpenStudy (lgg23):

i got it. the answer is 1/128

OpenStudy (anonymous):

Good. Out of curiousity did you use a calculator or convert to base ten logs first?

OpenStudy (lgg23):

Since the problem can be translated to \[\log_{2} x^{9/7}=9\]. I changed it to exponential form and got \[2^{9} = x^{9/7}\]

OpenStudy (lgg23):

then i got rid of the exponents by raising them to the reciprocal of that 9/7 and got (2^-7)=x and simplified

OpenStudy (anonymous):

Very nice - I like your method more. :-)

OpenStudy (lgg23):

Thank you.

OpenStudy (anonymous):

Just as it should be solved, exploit those log rules!! ;)